A certain game has two stacks of 30 tiles with pictures on them. In the first stack of tiles, there are 10 dogs, 4 cats, 5 balls, and 11 horses. In the second stack of tiles, there are 3 flowers, 8 fish, 12 balls, 2 cats, and 5 horses. The top tile in each stack is chosen. Find each probabillity.
1. P(neither is a horse)
2. P(exactly one is a fish)

Answer by stanbon(75887)

A certain game has two stacks of 30 tiles with pictures on them.
In the first stack of tiles, there are 10 dogs, 4 cats, 5 balls, and 11 horses.
----------
In the second stack of tiles, there are 3 flowers, 8 fish, 12 balls, 2 cats, and 5 horses.
--------
The top tile in each stack is chosen. Find each probabillity.
----
The events (selecting) are independent.
-------
1. P(neither is a horse) = (19/30)(25/30) = 0.5278
----------------------------
2. P(exactly one is a fish) = (1)(8/30) = 8/30
====================================================
Cheers,
Stan H.